OpenStudy (anonymous):

verify the identity, tan(x+pi/2)=-cot x

Nnesha (nnesha):

\[\huge\rm tan(x + y ) = \frac{ \tan x + \tan y }{ 1- \tan x \times \tan y }\] use this formula

Nnesha (nnesha):

pi/2 = 90 degree what is cos and sin value at 90 degree ?

OpenStudy (anonymous):

I'm not really sure.

Nnesha (nnesha):

nah forget that's wrong

Nnesha (nnesha):

*wait 4 someone else*

OpenStudy (anonymous):

Sure, thanks.

OpenStudy (anonymous):

\[\tan(x+\pi/2)=\frac{ \sin (x+\pi/2) }{ \cos(x+\pi/2) }=\frac{ sinx \cos \pi/2+\cos x \sin \pi/2 }{ \cos x \cos \pi/2-sinx \sin \pi/2 }\]

OpenStudy (anonymous):

\[\sin \pi/2=1 , \cos \pi/2 =0\]

OpenStudy (anonymous):

That verifies the identity?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thanks.